OpenStudy (anonymous):

Jack spends 1 1/4 as long on his homework as Jill. Jill spent 6 1/4 on her homework How long did Jack spend?

OpenStudy (texaschic101):

1 1/4 x 6 1/4 5/4 x 25/4 = 125/16 = 7 13/16 I think....you might want to get a second opinion....the wording kinda confuses me :)